71/7 divided by 8/5 =
Answer by Maths68(1474)

71/7 divided by 8/5 = 71/7/8/5
71/7 divided by 8/5 = 71/7*5/8
71/7 divided by 8/5 = 355/56
71/7 divided by 8/5 = 6 19/56 or 6.3392